# VDGD: Mitigating LVLM Hallucinations in Cognitive Prompts by Bridging the Visual Perception Gap

### v1

@@ 7 Conclusion, Limitations and Future Work
As part of future work, we would like to work on the current limitations of our work, which include: (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
@@ Appendix A Appendix
    B: We provide a detailed explanation of computational complexity of VDGD algorithm described in section 6.
@@
    C: We describe the impact of the paper on society and research communities.
@@
    D: We provide the prompts used for object detection, description, rephrasing, noise removal and evaluations required for our experiments.
@@
    E: We provide a detailed description of all the datasets utilized in this paper.
@@
    F: We provide a detailed description of all the models we have experimented with.
@@
    G: We provide a detailed description of all the prior-art hallucination mitigation techniques used as baselines.
@@
    H: We provide details about:
    • H.1: We provide detailed explanation about Logit Space for VR and Cognitive Prompts.
    • H.2: We provide examples comparing generations from Greedy, Decoding, VCD and VDGD.
    • H.3: We provide examples of noisy samples in datasets.
    • H.4: We provide and in-depth explanation of all the details of Algorithm 1.
    • H.5: We explain the elbow method used for logit analysis.
    • H.6: We provide examples of rephrased prompts which do not require image input.
    • H.7: We provide details of computational resources used for this paper.
@@
    I: We provide additional results for all the experiments described in the paper.
Appendix B VDGD Computational Complexity
@@
Given a LVLM with a vocabulary size $V$, an input instruction $I$ with a description of length $M$ tokens, at each decoding step the complexity for equation (2) is $O(V)$ and let the number of plausible tokens after this step be $K$, the complexity for equation (3) is $O(VMK)$. The total computational complexity of VDGD at each decoding step is $O(V)+O(VMK)\approx O(VMK)$.
Appendix C Broader Impact
@@
This paper addresses the broader impact of reducing hallucinations in Large Vision-Language Models (LVLMs), which are increasingly utilized in diverse applications such as automated content generation, assistive technologies, and decision-making systems. By focusing on the phenomenon of hallucination, particularly in responses to cognitive prompts, our work has the potential to enhance the reliability and safety of LVLMs in practical scenarios. The proposed Visual Description Grounded Decoding (VDGD) method, being training-free, offers an accessible and scalable solution that can be readily implemented across various models to mitigate risks associated with inaccurate or misleading outputs. Furthermore, the introduction of the VaLLu benchmark contributes to the field by providing a tool for comprehensive evaluation of cognitive capabilities in LVLMs, promoting transparency and fostering further research in the area. Collectively, these contributions aim to improve user trust in AI systems and encourage responsible AI development that prioritizes factual accuracy and understanding.
Appendix D Prompts
@@
For image description benchmarks we employ the prompt illustrated in Fig. 11. For non-image description benchmarks, we modify this prompt slightly (we modify the part of the prompt for extracting hallucinated phrases) and we illustrate the same in Fig. 15. The prompt used to generate image descriptions for VDGD is illustrated in Fig. 14. Correctness in the prompts refers to Factuality in all tables in Section I. We name it so after an ablation study on the terminology that works better for prompting the judge (GPT-4). All prompt evaluations on GPT-4 are done with a temperature of 0.7.
@@
Figure 15: Prompt used to evaluate generations from benchmarks like VaLLu, MMMU, MathVista, and SynthDoG using GPT-4-Turbo for the results in Table 2.
@@
Appendix E Dataset Details
@@
AMBER AMBER [29] is a benchmark for evaluating hallucinations in both the generative task and discriminative task of MLLMs. It provides comprehensive coverage of evaluations for various types of hallucination, including existence, attribute, and relation. Licensed under Apache-2.0 License.
@@
MMMU MMMU [39] benchmark evaluates multimodal models on multi-discipline tasks and demands college-level subject knowledge and deliberate reasoning. It consists of 11.5k questions, spanning 30 subjects and 183 subfields, focusing on advanced perception and reasoning on domain-specific knowledge. Licensed under Apache-2.0 License.
@@
MMC MMC [18] is a human-annotated benchmark comprising 9 distinct tasks to evaluate LVLM’s reasoning capabilities to comprehend visual charts. It comprises tasks such as chart information extraction, chart reasoning, contextual chart understanding, etc. It also offers two evaluation methods: GPT4 free-format general ability evaluation and multiple-choice QA format chart understanding. Licensed for research purposes.
@@
MathVista MathVista [23] is a robust mathematical reasoning evaluation benchmark that consists of challenging tasks that require fine-grained, deep visual recognition and compositional reasoning. It consists of 6141 examples derived from 31 multimodal datasets involving mathematics. Licensed under CC-BY-SA-4.0.
@@
MATH-Vision MATH-Vision [30] dataset is a collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions. The dataset spans 16 distinct mathematical disciplines and graded across 5 levels of difficulty to evaluate the mathematical reasoning abilities of Multimodal Large Language Models(MLLM). Licensed under MIT License.
@@
HallusionBench HallusionBench [9] is a comprehensive benchmark designed for the evaluation of image-context reasoning. The benchmark presents significant challenges to advanced large visual-language models (LVLMs), such as GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA- 1.5, by emphasizing nuanced understanding and interpretation of visual data. The benchmark comprises 346 images paired with 1129 questions, all meticulously crafted by human experts. Licensed under BSD-3-Clause.
@@
MME Multimodal large language model Evaluation Benchmark (MME) [6] measures perception and cognition abilities on a total of 2 tasks(Perception and Congition) and 14 subtasks(Existence, Count, Position, Color, Poster, Celebrity, Scene, Landmark, Artwork, OCR, Commonsense Reasoning, Numerical Calculation, Text Translation and Code Reasoning). To prevent data leakage from use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed.
@@
Oven Open-domain Visual Entity Recognition (OVEN) [12] is a task where a model needs to link an image onto a Wikipedia entity with respect to a text query. OVEN challenges models to select among six million possible Wikipedia entities, making it a general visual recognition benchmark with the largest number of labels. Licensed under MIT License.
@@
SynthDoG SynthDoG [14] is created by generating synthetic data which consists of various components: background, document, text, and layout. The background image is sampled from ImageNet, and the document is sampled from the collected paper photos. The text is sampled from Wikipedia. Layout is constructed using randomly stacks grids. The english variant of the dataset consists of 65.5k training and 500 validation entries. Licensed under MIT License.
Appendix F Model Details
@@
LLaVa-v1. LLaVa-v1 [22] [^1] https://huggingface.co/liuhaotian/llava-llama-2-7b-chat-lightning-lora-preview is an open-source LVLM trained by fine-tuning LLaMA/Vicuna (13B) on GPT-generated multimodal instruction-following data. It is initially trained on 558K filtered image-text pairs from LAION/CC/SBU and then, finetuned on 80K GPT-generated multimodal instruction-following data. Licensed under CC BY-SA 4.0 DEED.
@@
LLaVa-1.5. LLaVa-1.5 [20] [^2] https://huggingface.co/liuhaotian/llava-v1.5-7b builds on LLaVA-v1’s architecture by replacing linear projection design with a two-layer MLP. It also includes additional academic-task-oriented VQA datasets for VQA, OCR, and region-level perception, to enhance the model’s multimodal capabilities. Licensed under CC BY-SA 4.0 DEED.
@@
LLaVa-1.6. LLaVa-1.6 [21] [^3] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b re-uses the pretrained connector of LLaVA-1.5 and is instruction tuned on 158K GPT-generated multimodal data, 500K academic-task-oriented VQA data, 50K GPT-4V data and 40K ShareGPT data. LLaVa-1.6 has improved reasoning, OCR, and world knowledge. Licensed under CC BY-SA 4.0 DEED.
@@
mPLUG-Owl2. mPLUG-Owl2 [36] [^4] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b is the first Multi-modal Large Language Model that performs well in both pure-text and multi-modal scenarios. mPLUG-Owl2 is pretrained 400 million image-text pairs from: Conceptual Captions (CC3M/CC12M), COCO, Laio-en, COYO, DataComp and for instruction tuning 5 types of data are used: image captioning, mage question answering, region-aware QA, multi-modal instruct data and text-only instruct. Licensed under MIT License.
@@
InternLM-X. InternLM-X [40] [^5] https://huggingface.co/internlm/internlm-xcomposer2-vl-7b is a vision-language large model based on InternLM2 [3] for advanced text-image comprehension and composition. InternLM-X is pre-trained on 1.1 billion images alongside 77.7 billion text tokens, including both public datasets and in-house concept data followed by supervised fine-tuning using multi-task training and instruction tuning. Licensed under Apache-2.0 License.
@@
CogVLM. CogVLM [31] [^6] https://huggingface.co/THUDM/cogvlm-chat-hf is a powerful open-source visual language model. CogVLM-17B has 10 billion vision parameters and 7 billion language parameters. CogVLM is pre-trained using 1.5B text-image pairs and instruction tuned on various visual question-answering datasets. CogVLM achieves state-of-the-art performance on 10 classic cross-modal benchmarks. Licensed under Apache-2.0 License.
@@
Llama-2-7b Llama-2-7b [28] [^7] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf is an open source fine-tuned generative text model with 7 billion parameters. It is an auto-regressive language model that uses an optimized transformer architecture. The fine-tuned version use supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) for helpfulness and safety. Licensed under Llama 2 community license agreement.
@@
Vicuna-7b-v1.5 Vicuna-7b-v1.5 [45] [^8] https://huggingface.co/lmsys/vicuna-7b-v1.5 is a 7b parameter open-sourced model and fine-tuned from Llama 2 with supervised instruction fine-tuning and is an auto-regressive language model based on the transformer architecture. The training data is around 125K conversations from ShareGPT. Licensed under Apache-2.0 License.
@@
InternLM2-7B InternLM-7B [3] [^9] https://huggingface.co/internlm/internlm2-chat-7b InternLM2 is a open-source 7 billion parameter chat model with a 200k context window and achieves good performance in the“Needle-in-a-Haystack” test(on long-context tasks like LongBench and L-Eval) and is trained using Conditional Online RLHF (COOL RLHF). Licensed under Apache-2.0 License.
Appendix G Baseline Details
@@
VCD. Visual Contrastive Decoding (VCD) [26] is a training-free method that contrasts output distributions of original/unaltered inputs and distorted visual inputs. VCD hypothesizes statistical bias and unimodal priors as the two essential causes of object hallucinations and reduces the dependence on them. This results in the generated content being closely grounded to visual inputs resulting more accurate outputs. Licensed under Apache-2.0 License.
@@
OPERA. Over-trust Penalty and a Retrospection-Allocation (OPERA) [13] is a training-free approach that hypothesizes that hallucinations are closely tied to the knowledge aggregation patterns in the self-attention matrix and that generation of new tokens by models is majorly dependent on a few summary tokens(over-trust issue) but not all of the previous tokens. This methodology introduces a penalty term on the model logits during beam-search decoding to mitigate the over-trust issue and implements a rollback strategy to retrospect the presence of summary tokens. Licensed under MIT License.
@@
Woodpecker. Woodpecker [38] is a training-free method that identifies and corrects hallucinations in the generated text. Unlike other training-free methodologies, Woodpecker works after the model generation process. This method works by creating a visual knowledge base of Question-Answer(QA) pairs which contain object-level and attribute-level claims about the input image followed by an LLM correcting hallucinations.
@@
LRV. LRV [17], or Large-scale Robust Visual Instruction tuning, proposes a dataset of visual instructions generated by GPT4, covering 16 vision-and-language tasks with open-ended instructions and answers. Each instance has positive instruction samples and negative instructions for more robust visual instruction tuning. As a result, a model trained on this dataset proves to be robust to hallucinations. Licensed under BSD-3-Clause license.
@@
LURE. LVLM Hallucination Revisor (LURE) [47] attributes object hallucination to three factors: co-occurrence, uncertainty, and object position, and develops an object hallucination revisor which converts potential hallucinations into accurate outputs. This method first creates a hallucination dataset by making modifications to the original data and trains the object hallucination revisor on this dataset. This method post-hoc rectifies object hallucinations in LVLMs by reconstructing less hallucinatory descriptions.
Appendix H Additional Details
H.1 Logit Space for VR and Cognitive Prompts
@@
H.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
In this subsection, we illustrate several instances (figures 17, 20, 21, 18, 19) from VaLLu and compare their responses for LLaVa-1.5 with Greedy, VCD, and VDGD Decoding.
@@ H.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
H.3 Noisy Examples in Existing Datasets
To maintain a high quality of the benchmark and eliminate noisy examples we performed a manual review of the existing datasets. This was crucial to ensure that the data used in our paper was reliable. It was also essential for the validity of our research findings. We removed questions that did not have a definite ground truth answer. We also made sure that the questions were answerable using a combination of general knowledge and reasoning. This filtering was done by the paper authors and has been approved by our institution’s Institutional Review Board (IRB). We share a few examples (figures: 22, 23, 24, 25, 26, 27, 28, 29) of noise in the existing datasets below:
@@ H.3 Noisy Examples in Existing Datasets
H.4 Additional Details on Algorithm 1

### v2

@@ 6 Conclusion, Limitations and Future Work
As part of future work, we would like to work on the current limitations of our work, which include (1) Error accumulation in VDGD that comes from inaccurate image descriptions in the prefix, (2) the requirement of prompting the LVLM twice, (3) Analysis of our findings with better alignment techniques and methods like RAG.
7 Reproducibility Statement
We provide our code here: https://anonymous.4open.science/r/VDGD-1E04/. All codes will be open-sourced upon paper acceptance, including all checkpoints. All experimental details, including training parameters and hyper-parameters, are provided in Section 5.3. The Appendix has more comprehensive details about our experimental setup.
@@ Appendix A Appendix
    B: We illustrate failure cases for VDGD.
@@
    C: We provide results for ablation of VDGDs key components.
@@
    D: We provide correlation between GPT scores and actual metrics.
@@
    E: We provide a detailed explanation of computational complexity of VDGD algorithm described in section 5.
@@
    F: We describe the impact of the paper on society and research communities.
@@
    G: We provide the prompts used for object detection, description, rephrasing, noise removal and evaluations required for our experiments.
@@
    H: We provide a detailed description of all the datasets utilized in this paper.
@@
    I: We provide a detailed description of all the models we have experimented with.
  9. 9.
    J: We provide a detailed description of all the prior-art hallucination mitigation techniques used as baselines.
  10. 10.
    K: We provide details about:
    • K.1: We provide detailed explanation about Logit Space for VR and Cognitive Prompts.
    • K.2: We provide examples comparing generations from Greedy, Decoding, VCD and VDGD.
    • K.3: We provide examples of noisy samples in datasets.
    • K.4: We provide and in-depth explanation of all the details of Algorithm 1.
    • K.6: We explain the elbow method used for logit analysis.
    • K.7: We provide examples of rephrased prompts which do not require image input.
    • K.8: We provide details of computational resources used for this paper.
  11. 11.
    L: We provide additional results for all the experiments described in the paper.
Appendix B VDGD Failure Cases
Fig. 11 illustrates 2 failure cases of VDGD. We show that VDGD can lead to incorrect responses due to an incorrectly generated image description. However, we also argue that as LVLMs get better at describing images, the performance of VDGD decoding will improve.
[…]
[width=]failure_cases.pdf
Figure 11: Illustration of failure cases of VDGD.
Appendix C VDGD Ablation Results
To demonstrate the effectiveness of VDGD, we conduct ablation studies on its key components: (i) VDGD (- VCD): Image descriptions are generated without using Visual Contrastive Decoding (VCD), (ii) VDGD (- KLD): KLD-based sampling is removed, and only the description is appended, (iii) VDGD (- Image): VDGD is applied without the original image input; only the image description is appended to the prompt. Results are presented on the VDGD benchmark in Table 3. Removing KL Divergence leads to a substantial performance drop. We notice a similar phenomenon by removing the input image with the steepest drop. This shows that the input image is still essential for the model.
[…]
Table 3: Ablation results for VDGD. VDGD sees a drop in performance when the KLD decoding objective is removed and a major drop when no input image is provided.
Appendix D Correlation between GPT scores and actual metrics
Background of Experts in Human Evaluation. For the expert human evaluation of model generations, we engaged a group of 5 PhD students with at least four years of research experience in computer vision. These experts were tasked with evaluating the outputs of the models for accuracy, relevance, and factual correctness. To ensure thorough evaluation, they were provided with internet access to verify factual information when necessary. The evaluation process was conducted in an anonymized format to maintain objectivity, with the experts working independently and without prior knowledge of the model specifics. We selected 500 samples or the size of the benchmark, whichever was lower.
Appendix E VDGD Computational Complexity
@@
Given an LVLM with a vocabulary size $V$, an input instruction $I$ with a description of length $M$ tokens, at each decoding step, the complexity for equation (2) is $O(V)$ and let the number of plausible tokens after this step be $K$, the complexity for equation (3) is $O(VMK)$. The total computational complexity of VDGD at each decoding step is $O(V)+O(VMK)\approx O(VMK)$.
Appendix F Broader Impact
@@
This paper addresses the broader impact of reducing hallucinations in Large Vision-Language Models (LVLMs), which are increasingly utilized in diverse applications such as automated content generation, assistive technologies, and decision-making systems. By focusing on the phenomenon of hallucination, particularly in responses to cognitive prompts, our work has the potential to enhance the reliability and safety of LVLMs in practical scenarios. The proposed Visual Description Grounded Decoding (VDGD) method, being training-free, offers an accessible and scalable solution that can be readily implemented across various models to mitigate risks associated with inaccurate or misleading outputs. Furthermore, the introduction of the VaLLu benchmark contributes to the field by providing a tool for the comprehensive evaluation of cognitive capabilities in LVLMs, promoting transparency and fostering further research in the area. Collectively, these contributions aim to improve user trust in AI systems and encourage responsible AI development that prioritizes factual accuracy and understanding.
Appendix G Prompts
@@
For image description benchmarks, we employ the prompt illustrated in Fig. 12. For non-image description benchmarks, we modify this prompt slightly (we modify the part of the prompt for extracting hallucinated phrases), and we illustrate the same in Fig. 16. The prompt used to generate image descriptions for VDGD is illustrated in Fig. 15. Correctness in the prompts refers to Factuality in all tables in Section L. We name it so after an ablation study on the terminology that works better for prompting the judge (GPT-4). All prompt evaluations on GPT-4 are done with a temperature of 0.7.
[…]
[width=]evalutaion_prompt.pdf
@@
[…]
[width=]maths_eval_prompt.pdf
Figure 16: Prompt used to evaluate generations from cognitive benchmarks in Table 9.
[…]
[width=]prompt_rebut.pdf
Figure 17: GPT evaluation prompt for evaluating informativeness of image descriptions generated by LVLMs.
[…]
[width=]text-only-eval.pdf
@@
Appendix H Dataset Details
@@
AMBER AMBER (Wang et al., 2023a) is a benchmark for evaluating hallucinations in both the generative task and discriminative task of MLLMs. It provides comprehensive coverage of evaluations for various types of hallucination, including existence, attribute, and relation. Licensed under Apache-2.0 License.
@@
MMMU MMMU (Yue et al., 2024) benchmark evaluates multimodal models on multi-discipline tasks and demands college-level subject knowledge and deliberate reasoning. It consists of 11.5k questions, spanning 30 subjects and 183 subfields, focusing on advanced perception and reasoning on domain-specific knowledge. Licensed under Apache-2.0 License.
@@
MMC MMC (Liu et al., 2023a) is a human-annotated benchmark comprising 9 distinct tasks to evaluate LVLM’s reasoning capabilities to comprehend visual charts. It comprises tasks such as chart information extraction, chart reasoning, contextual chart understanding, etc. It also offers two evaluation methods: GPT4 free-format general ability evaluation and multiple-choice QA format chart understanding. Licensed for research purposes.
@@
MathVista MathVista (Lu et al., 2023) is a robust mathematical reasoning evaluation benchmark that consists of challenging tasks that require fine-grained, deep visual recognition and compositional reasoning. It consists of 6141 examples derived from 31 multimodal datasets involving mathematics. Licensed under CC-BY-SA-4.0.
@@
MATH-Vision MATH-Vision (Wang et al., 2024a) dataset is a collection of 3,040 high-quality mathematical problems with visual contexts sourced from real math competitions. The dataset spans 16 distinct mathematical disciplines and is graded across 5 levels of difficulty to evaluate the mathematical reasoning abilities of Multimodal Large Language Models(MLLM). Licensed under MIT License.
@@
HallusionBench HallusionBench (Guan et al., 2023) is a comprehensive benchmark designed for the evaluation of image-context reasoning. The benchmark presents significant challenges to advanced large visual-language models (LVLMs), such as GPT-4V(ision), Gemini Pro Vision, Claude 3, and LLaVA- 1.5, by emphasizing nuanced understanding and interpretation of visual data. The benchmark comprises 346 images paired with 1129 questions, all meticulously crafted by human experts. Licensed under BSD-3-Clause.
LlavaBench LlavaBench (Liu et al., 2023c) is a challenging evaluation suite designed to assess visual-language alignment and instruction-following capabilities of large visual-language models. It consists of two benchmarks: LLaVA-Bench (COCO), which includes 30 randomly selected images from COCO-Val-2014 paired with 90 questions covering conversational, descriptive, and complex reasoning tasks, and LLaVA-Bench (In-the-Wild), featuring 24 diverse images from various domains, including memes, paintings, and sketches, with 60 questions.
MM-Vet MM-Vet (Yu et al., 2023) is an evaluation benchmark designed to assess large multimodal models on complex multimodal tasks. It defines six core vision-language capabilities—Recognition, Knowledge, OCR, Spatial Awareness, Language Generation, and Math—and examines their integration across 16 emergent tasks. MM-Vet employs an LLM-based evaluator for open-ended outputs, providing unified scoring across diverse question types and answer styles. The benchmark offers insights into the capabilities of different large multimodal model paradigms and models beyond simple performance ranking.
@@
MME Multimodal large language model Evaluation Benchmark (MME) (Fu et al., 2023) measures perception and cognition abilities on a total of 2 tasks(Perception and Congition) and 14 subtasks(Existence, Count, Position, Color, Poster, Celebrity, Scene, Landmark, Artwork, OCR, Commonsense Reasoning, Numerical Calculation, Text Translation and Code Reasoning). To prevent data leakage from use of public datasets for evaluation, the annotations of instruction-answer pairs are all manually designed.
@@
Oven Open-domain Visual Entity Recognition (OVEN) (Hu et al., 2023) is a task where a model needs to link an image onto a Wikipedia entity with respect to a text query. OVEN challenges models to select among six million possible Wikipedia entities, making it a general visual recognition benchmark with the largest number of labels. Licensed under MIT License.
@@
SynthDoG SynthDoG (Kim et al., 2022) is created by generating synthetic data which consists of various components: background, document, text, and layout. The background image is sampled from ImageNet, and the document is sampled from the collected paper photos. The text is sampled from Wikipedia. Layout is constructed using randomly stacks grids. The english variant of the dataset consists of 65.5k training and 500 validation entries. Licensed under MIT License.
Appendix I Model Details
@@
LLaVa-v1. LLaVa-v1 (Liu et al., 2023c) [^3] https://huggingface.co/liuhaotian/llava-llama-2-7b-chat-lightning-lora-preview is an open-source LVLM trained by fine-tuning LLaMA/Vicuna (13B) on GPT-generated multimodal instruction-following data. It is initially trained on 558K filtered image-text pairs from LAION/CC/SBU and then, finetuned on 80K GPT-generated multimodal instruction-following data. Licensed under CC BY-SA 4.0 DEED.
@@
LLaVa-1.5. LLaVa-1.5 (Liu et al., 2023b) [^4] https://huggingface.co/liuhaotian/llava-v1.5-7b builds on LLaVA-v1’s architecture by replacing linear projection design with a two-layer MLP. It also includes additional academic-task-oriented VQA datasets for VQA, OCR, and region-level perception, to enhance the model’s multimodal capabilities. Licensed under CC BY-SA 4.0 DEED.
@@
LLaVa-1.6. LLaVa-1.6 (Liu et al., 2024c) [^5] https://huggingface.co/liuhaotian/llava-v1.6-vicuna-7b re-uses the pretrained connector of LLaVA-1.5 and is instruction tuned on 158K GPT-generated multimodal data, 500K academic-task-oriented VQA data, 50K GPT-4V data and 40K ShareGPT data. LLaVa-1.6 has improved reasoning, OCR, and world knowledge. Licensed under CC BY-SA 4.0 DEED.
@@
mPLUG-Owl2. mPLUG-Owl2 (Ye et al., 2023) [^6] https://huggingface.co/MAGAer13/mplug-owl2-llama2-7b is the first Multi-modal Large Language Model that performs well in both pure-text and multi-modal scenarios. mPLUG-Owl2 is pretrained 400 million image-text pairs from: Conceptual Captions (CC3M/CC12M), COCO, Laio-en, COYO, DataComp and for instruction tuning 5 types of data are used: image captioning, mage question answering, region-aware QA, multi-modal instruct data and text-only instruct. Licensed under MIT License.
@@
InternLM-X. InternLM-X (Zhang et al., 2023) [^7] https://huggingface.co/internlm/internlm-xcomposer2-vl-7b is a vision-language large model based on InternLM2 (Cai et al., 2024) for advanced text-image comprehension and composition. InternLM-X is pre-trained on 1.1 billion images alongside 77.7 billion text tokens, including both public datasets and in-house concept data followed by supervised fine-tuning using multi-task training and instruction tuning. Licensed under Apache-2.0 License.
@@
CogVLM. CogVLM (Wang et al., 2023b) [^8] https://huggingface.co/THUDM/cogvlm-chat-hf is a powerful open-source visual language model. CogVLM-17B has 10 billion vision parameters and 7 billion language parameters. CogVLM is pre-trained using 1.5B text-image pairs and instruction tuned on various visual question-answering datasets. CogVLM achieves state-of-the-art performance on 10 classic cross-modal benchmarks. Licensed under Apache-2.0 License.
@@
Llama-2-7b. Llama-2-7b (Touvron et al., 2023) [^9] https://huggingface.co/meta-llama/Llama-2-7b-chat-hf is an open source fine-tuned generative text model with 7 billion parameters. It is an auto-regressive language model that uses an optimized transformer architecture. The fine-tuned version use supervised fine-tuning (SFT) and reinforcement learning with human feedback (RLHF) for helpfulness and safety. Licensed under Llama 2 community license agreement.
@@
Vicuna-7b-v1.5. Vicuna-7b-v1.5 (Zheng et al., 2023) [^10] https://huggingface.co/lmsys/vicuna-7b-v1.5 is a 7b parameter open-sourced model and fine-tuned from Llama 2 with supervised instruction fine-tuning and is an auto-regressive language model based on the transformer architecture. The training data is around 125K conversations from ShareGPT. Licensed under Apache-2.0 License.
@@
InternLM2-7B. InternLM-7B (Cai et al., 2024) [^11] https://huggingface.co/internlm/internlm2-chat-7b InternLM2 is a open-source 7 billion parameter chat model with a 200k context window and achieves good performance in the“Needle-in-a-Haystack” test(on long-context tasks like LongBench and L-Eval) and is trained using Conditional Online RLHF (COOL RLHF). Licensed under Apache-2.0 License.
CogVLM2. CogVLM2 (Hong et al., 2024) [^12] https://huggingface.co/THUDM/cogvlm2-llama3-chat-19B CogVLM2 is a new generation of visual language models based on Meta-LLaMA-3-8B-Instruct, offering significant improvements over previous versions. It supports 8K content length, image resolutions up to 1344x1344, and excels in both English and Chinese. CogVLM2 has achieved state-of-the-art performance on benchmarks like TextVQA, DocVQA, and MM-Vet. The open-source models, including CogVLM2 and CogVLM2-Video, are available on GitHub, contributing to advancements in image and video understanding. Licensed under Apache-2.0 License.
Qwen2-VL. Qwen2-VL (Wang et al., 2024b) [^13] https://huggingface.co/Qwen/Qwen2-VL-7B-Instruct Qwen2-VL is the latest iteration of the Qwen-VL model, featuring state-of-the-art performance in image and video understanding tasks. It excels in processing images of various resolutions and aspect ratios and supports long-form video understanding for tasks like question answering and content creation. Key architectural updates include Naive Dynamic Resolution and Multimodal Rotary Position Embedding (M-ROPE), improving its multimodal processing capabilities. Available in 2, 7, and 72 billion parameter variants. Licensed under Apache-2.0 License.
Appendix J Baseline Details
@@
VCD. Visual Contrastive Decoding (VCD) (Sicong Leng et al., 2023) is a training-free method that contrasts output distributions of original/unaltered inputs and distorted visual inputs. VCD hypothesizes statistical bias and unimodal priors as the two essential causes of object hallucinations and reduces the dependence on them. This results in the generated content being closely grounded to visual inputs resulting more accurate outputs. Licensed under Apache-2.0 License.
@@
OPERA. Over-trust Penalty and a Retrospection-Allocation (OPERA) (Huang et al., 2023b) is a training-free approach that hypothesizes that hallucinations are closely tied to the knowledge aggregation patterns in the self-attention matrix and that generation of new tokens by models is majorly dependent on a few summary tokens(over-trust issue) but not all of the previous tokens. This methodology introduces a penalty term on the model logits during beam-search decoding to mitigate the over-trust issue and implements a rollback strategy to retrospect the presence of summary tokens. Licensed under MIT License.
@@
Woodpecker. Woodpecker (Yin et al., 2023b) is a training-free method that identifies and corrects hallucinations in the generated text. Unlike other training-free methodologies, Woodpecker works after the model generation process. This method works by creating a visual knowledge base of Question-Answer(QA) pairs which contain object-level and attribute-level claims about the input image followed by an LLM correcting hallucinations.
@@
LRV. LRV (Liu et al., 2024a), or Large-scale Robust Visual Instruction tuning, proposes a dataset of visual instructions generated by GPT4, covering 16 vision-and-language tasks with open-ended instructions and answers. Each instance has positive instruction samples and negative instructions for more robust visual instruction tuning. As a result, a model trained on this dataset proves to be robust to hallucinations. Licensed under BSD-3-Clause license.
@@
LURE. LVLM Hallucination Revisor (LURE) (Zhou et al., 2023b) attributes object hallucination to three factors: co-occurrence, uncertainty, and object position, and develops an object hallucination revisor which converts potential hallucinations into accurate outputs. This method first creates a hallucination dataset by making modifications to the original data and trains the object hallucination revisor on this dataset. This method post-hoc rectifies object hallucinations in LVLMs by reconstructing less hallucinatory descriptions.
PAI. Pay Attention to Image (PAI) (Liu et al., 2024d) is a training-free algorithm designed to reduce hallucinations in LVLMs. PAI intervenes during the inference process to make it more image-centric by amplifying attention to image tokens in the self-attention layers of LVLMs. By adjusting attention weights and subtracting logits of text-only inputs from multi-modal inputs, PAI mitigates the issue of “text inertia,” where outputs are overly influenced by context text. This method enhances image representation during generation, leading to more accurate, visually grounded outputs without requiring additional training or external tools.
HALC. HALC (Chen et al., 2024a) is a novel decoding algorithm designed to mitigate object hallucinations (OH) in LVLMs. It integrates a robust auto-focal grounding mechanism to correct hallucinated tokens locally and a specialized beam search algorithm to reduce OH globally while preserving text generation quality. HALC can be easily integrated into any LVLM without additional training, addressing all three types of OH—existence, attribute, and relationship. Extensive experiments show that HALC outperforms state-of-the-art methods across multiple benchmarks.
Appendix K Additional Details
K.1 Logit Space for VR and Cognitive Prompts
@@
K.2 Qualitative Comparison of Greedy vs VCD vs VDGD Decoding
In this subsection, we illustrate several instances (figures 20, 22, 23, 21, 24) from VaLLu and compare their responses for LLaVa-1.5 with Greedy, VCD, and VDGD Decoding.
[…]
[width=]qual_with_desc.pdf
Figure 19: Qualitative example for VaLLu benchmark also showing the descriptions.
[…]
[width=]qual_app_img_1.pdf
@@
K.3 Noisy Examples in Existing Datasets
To maintain a high quality of the benchmark and eliminate noisy examples, we performed a manual review of the existing datasets. This was crucial to ensure that the data used in our paper was reliable. It was also essential for the validity of our research findings. We removed questions that did not have a definite ground truth answer. We also made sure that the questions were answerable using a combination of general knowledge and reasoning. This filtering was done by the paper authors and has been approved by our institution’s Institutional Review Board (IRB). We share a few examples (figures: 25, 26, 27, 28, 29, 30, 31, 32) of noise in the existing datasets below:
[…]
[width=]mmmu_noisy_img_1.pdf
@@
K.4 Additional Details on Algorithm 1